Which is the standard form of the equation where the point (-2,4) is on a circle whose center is the origin?
Answer by stanbon(75887)

Which is the standard form of the equation where the point (-2,4) is on a circle whose center is the origin?
---
radius equals the distance from (-2,4) to the origin.
---
r^2 2^2 + 4^2 = 20
---------------------------
r^2 = 20
---
Equation:
(x+2)^2 + (y-4)^2 = 20
==============================
Cheers,
Stan H.